# Federated nnU-Net for Privacy-Preserving Medical Image Segmentation

### v1

@@ 3.3.2 ACOUSLIC-AI Sierra Leone:
For the external validation, we utilize the publicly available samples from ACOUSLIC-AI [^5] https://acouslic-ai.grand-challenge.org/datasets/ Challenge dataset [30]. We use a collection of 178 cases collected in Sierra Leone from three Public Health Units (PHUs). With a resolution of 744×562 pixels, the images from Sierra Leone sit between the low-end devices used in Algeria and Malawi and the high-end scanners employed in Spain. This intermediate imaging quality makes the dataset an ideal candidate for assessing model generalizability across varying image qualities and clinical environments, ensuring robust validation of AI performance.
4.  Experiments
4.1.  Experimental setups
The training process is conducted across four different setups:
@@ 4.1.  Experimental setups
  • Single-center Training:  Each node independently trains its own nnU-Net model using only its local dataset. This setup serves as a baseline, reflecting the performance of a node/institution without participation in the federation.
@@
  • Centralized Training:   All data is aggregated into a single dataset, which is used to train a centralized model. This setup represents the theoretical upper bound of performance, though it is typically infeasible in real-world scenarios due to privacy constraints.
@@
  • Federated Fingerprint Training:   Utilizing our FFE approach (2), the local dataset fingerprints from all nodes are aggregated at the central node into one federated fingerprint and distributed back to each node. As a result, each training site obtains identical network architecture during the nnU-Net’s self-configuration process. This approach approximates the nnU-Net’s behavior in the centralized training setup.
@@
  • Asymmetric Aggregation Training:   Implemented using our AsymFedAvg strategy, allowing nodes to maintain their original nnU-Net configured architectures while sharing compatible parts of their networks.
@@
For our experimentation, we maintain nnU-Net’s default configuration parameters as described in [7]. We test the 2D U-Net architecture on breast MRI and fetal ultrasound data and the 3D architecture on a cardiac MRI use case, consistently employing the default training length of 1000 epochs and using the last model checkpoint for evaluation across all federated nodes. Following other recent works on medical image segmentation [7, 31, 32, 33], we evaluate the models with Dice Similarity Coefficient (DSC) and 95% Hausdorff distance (HD95). For internal validation, we report results as the average of a 5-fold cross-validation, while for external testing, we follow the original nnU-Net approach by generating ensemble predictions from all five folds. Moreover, each experiment uses identical data splits—for example, the n-th fold in the centralized setup corresponds directly to the set of n-th folds from centers used in the local setup. By adhering to these protocols, we try to minimize the influence of other factors on the final results, ensuring that any observed performance variations are primarily due to our federated learning adaptations rather than modifications to the base framework.
4.2.  Results
@@
Table 8: Quantitative results for multi-center MAMA-MIA and EuCanImage datasets. HD95 values reported in mm. Best results for privacy-preserving methods are in bold.
@@
Single-center 0.612 17.60 0.739 19.03 0.616 33.56 0.551 23.24 0.579 35.64 0.636 31.72 0.544 46.30
AsymFedAvg 0.634 16.69 0.739 18.15 0.621 28.70 0.620 21.79 0.557 51.43 0.637 32.43 0.551 49.10
FFE 0.644 16.20 0.742 17.48 0.653 39.54 0.660 17.49 0.582 28.83 0.645 33.52 0.565 40.41
Centralized 0.654 17.28 0.750 19.84 0.657 58.99 0.675 17.35 0.582 52.44 0.644 41.17 0.569 57.98
Table 9: Quantitative results for multi-center M&Ms and ACDC datasets. HD95 values reported in mm. Best results for privacy-preserving methods are in bold.
@@
Single-center 0.908 4.27 0.912 3.62 0.902 6.56 0.896 4.31 0.900 4.34 0.887 4.79
AsymFedAvg 0.909 4.20 0.913 3.70 0.907 4.23 0.897 4.29 0.902 4.07 0.889 4.58
FFE 0.911 4.05 0.914 3.67 0.922 3.30 0.904 3.98 0.904 3.87 0.897 3.87
Centralized 0.910 4.20 0.914 3.53 0.923 3.26 0.905 3.87 0.907 3.81 0.898 3.89
@@
Method Algeria Egypt Malawi Spain external Sierra Leone
@@
Single-center 0.952 11.87 0.931 24.21 0.926 26.83 0.942 29.82 0.831 53.20
AsymFedAvg 0.951 11.16 0.941 18.29 0.939 16.20 0.944 25.42 0.851 43.44
FFE 0.955 10.23 0.954 14.56 0.943 15.64 0.938 47.58 0.809 46.57
Centralized 0.954 10.23 0.953 14.50 0.943 15.55 0.932 55.76 0.804 54.34
(a) MAMA-MIA cross-center validation matrix
(b) M&Ms cross-center validation matrix
(c) AIMIX cross-center validation matrix
@@
4.2.1 Breast MRI
Table 8 presents the segmentation performance for the breast lesion MAMA-MIA dataset. Both proposed methods, FFE and AsymFedAvg, outperform the local training of individual centers in terms of DSC and HD95 across all datasets. Specifically, FFE method achieves an average DSC improvement of 3.2% on ISPY1, 0.3% on ISPY2, 3.7% on DUKE, and a notable 10.9% on NACT. Additionally, FFE maintains DSC scores close to the centralized approach, while surpassing it in terms of HD95, effectively bridging the performance gap in a decentralized framework. Interestingly for DUKE, AsymFedAvg outperforms every other method in terms of HD95. Since the DUKE dataset comprises samples of substantially different spatial characteristics (pixel resolution approximately 2 times higher, see Table 4), the adaptive nature of the asymmetric aggregation approach allows this center to maintain its initial architecture, while benefiting from the other centers. In this case, the HD95 gets reduced in contrast to unified architecture methods (FFE and centralized setting). The cross-center validation matrix in Fig. 4 further exposes the dataset’s heterogeneity, where both ISPY2 and NACT trained models struggle to generalize well for the DUKE dataset, with the only exception of ISPY2, which has image resolution much closer to DUKE. Sample predictions presented in Fig. 5 show the strong improvement of segmentation performance for NACT dataset, which struggles to perform well with its locally trained nnU-Net model. For DUKE, HD95 improvement can be observed for AsymFedAvg model prediction. External validation results for the EuCanImage dataset show that federated training with the FFE method achieves DSC scores on par with the centralized model. While the segmentation performance for the KAUNO center is comparable to the results achieved within MAMA-MIA dataset, demonstrating good generalizability. The GUMED center exhibits a notable decrease in performance, which is most likely attributed to discrepancies in annotation protocols rather than the data itself. An example of such difference is illustrated in Fig. 7, where fine-grained model predictions fail to align with coarse ground truth regions provided by experts.
Figure 5: Sample predictions of each method for Breast Cancer datasets.
4.2.2 Cardiac MRI
For M&Ms, similarly to previous dataset, both proposed federated learning methods outperform the local, non-collaborative training setup across both tested metrics. The only exception being Clinica Sagrada Familia, for which the HD95 is lower in the local setup by a minimal margin. FFE method consistently achieves the highest scores in terms of DSC for both internal and external datasets, while the adaptive AsymFedAvg does not reflect substantial benefits from federated training. Most likely, it is caused by dominantly homogeneous nature of M&Ms dataset, in which all data centers share similar sample characteristics making a common architecture method more suitable. This effect can be observed as well in the cross-center evaluation of locally trained models (see Fig. 4) where each center managed to provide acceptable results for samples from other, unseen institutions. Moreover, it is important to note that the FFE method, despite the additional privacy requirement manages to provide almost identical performance to the centralized training scenario. Sample predictions for cardiac MRI dataset can be observed in Fig. 6. All tested methods provide consistent results for central slices while having more visible disparities in regions close to the heart’s apex (sample predictions from ACDC) and heart’s atria (sample predictions from Vall d’Hebron).
@@ 4.2.2 Cardiac MRI
Figure 6: Sample predictions of each method for cardiac MRI datasets.
4.2.3 Fetal Ultrasound
For fetal abdomen segmentation task, considerable differences can be observed between FFE and AsymFedAvg methods; see Table 10. Due to strong data heterogeneity, i.e. discrepancies in capture devices and image resolution between centers (Table 7), the models trained with data from Algeria, Egypt or Malawi do not generalize well for higher resolution samples from Spain; see Fig. 4c. In this case, having a unified training strategy for all centers does not prove to be the optimal solution. Allowing the largest center, i.e. Spain, to keep its original architecture allows to get a more appropriate model to the characteristics of the external center, i.e. Sierra Leone, while still benefiting from the federated training and contribution of other centers. This effect can be observed on sample predictions presented in Fig. 7. The single-center nnU-Net model often struggled to identify the right region for abdomen delineation, which can be especially observed in the external validation samples from Sierra Leone.
Figure 7: Sample predictions of each method for AIMIX dataset.
5.  Discussion
@@
Our experiments across three diverse imaging use cases - breast MRI (MAMA-MIA), cardiac MRI (M&Ms), and fetal ultrasound (AIMIX) - show that FednnU-Net can reliably match, or even surpass conventional centralized strategies. These results add to a growing body of evidence illustrating that distributed model optimization can achieve comparable accuracy while respecting the complexities of real-world clinical collaborations, such as data privacy and varied institution-specific protocols. Moreover, our findings underscore the critical need for methods that effectively handle data heterogeneity during federated model training.
Whilst other, federated state-of-the-art segmentation methods may achieve higher accuracy with highly customized solutions for specific tasks, the primary objective of this work is to provide an easy-to-use, privacy-preserving solution accessible to researchers with varying levels of expertise. In doing so, we prioritize the broader applicability and real-world viability of the framework over achieving the absolute highest segmentation performance.
The increasing emphasis on privacy-preserving, decentralized settings is rapidly becoming the standard in modern AI solutions — especially within healthcare, where patient data confidentiality is paramount. FednnU-Net aligns with this trend by enabling collaborative training without compromising data privacy. However, several aspects require further investigation:
@@
    Scalability and Client Diversity: Future studies should test FednnU-Net with a larger number of nodes to better mimic real-world deployments, where numerous and geographically dispersed institutions may challenge model convergence and communication.
@@
    Heterogeneous Computational Resources: Given that clients often have different processing capabilities, it is crucial to evaluate FednnU-Net under varied hardware constraints. Although our framework already supports such a scenario, this dimension was not explored in this current study.
@@
    Enhanced Privacy Mechanisms: Beyond our current privacy measures, integrating techniques such as differential privacy [34, 35] and secure multiparty computation (SMPC) could further reduce the risk of data leakage, even in aggregated updates.
In summary, although more accurate medical segmentation techniques exist, FednnU-Net fills a vital gap by balancing segmentation performance with ease-of-use and privacy preservation. As federated learning continues to mature and becomes an increasingly standard practice in healthcare, addressing these limitations through further evaluation and enhancements will be essential for broader adoption and improved robustness in clinical applications.
6.  Conclusion
In this study, we present FednnU-Net, a federated learning framework for 2D and 3D medical image segmentation that bridges the gap between privacy-sensitive setups and high-performance model training. Our framework introduces two methods: Federated Fingerprint Extraction (FFE) and Asymetric Federated Averaging (AsymFedAvg) that unlock the potential of distributed model training under privacy constraints. FFE enables construction of one, universal model architecture akin to centralized training, ensuring robustness across data centers. Alternatively, AsymFedAvg allows tailored model customization for each center to better align with local data characteristics. Our results demonstrate that both FFE and AsymFedAvg offer distinct performance benefits depending on the multi-center dataset characteristics, either achieving performance close to the centralized setting or surpassing it, particularly in highly diverse datasets. On top of that, both approaches preserve nnU-Net’s hallmark self-configuring capabilities, and due to our modular open-source implementation, they allow researchers to seamlessly translate their nnU-Net-derived centralized segmentation methods to the federated setting. Furthermore, we open source our framework and release data enhancing two public dataset collections to allow reproducibility of results. We hope that our contributions provide a robust foundation for further development and evaluation of medical image segmentation models in privacy-constrained settings.

### v2

@@ Methods
For the external validation, we utilize the publicly available samples from ACOUSLIC-AI Challenge dataset [44]. We use a collection of 178 cases collected in Sierra Leone from three Public Health Units (PHUs). With a resolution of 744×562 pixels, the images from Sierra Leone sit between the low-end devices used in Algeria and Malawi and the high-end scanners employed in Spain. This intermediate imaging quality makes the dataset an ideal candidate for assessing model generalizability across varying image qualities and clinical environments, ensuring robust validation of AI performance.
Experiments
Experimental setups
We adopt the vanilla nnU-Net architecture in both local and centralized training setups as the primary baselines for evaluating the performance of our proposed FednnU-Net framework. This choice is grounded in the fact that nnU-Net has established itself as a strong, standardized, and adaptive baseline for medical image segmentation tasks across a wide range of datasets [14]. By using the same framework and memory constraints in local (per-institution) and centralized (pooled data) settings, we ensure a fair and controlled comparison that isolates the effect of the federation process itself, without introducing variability due to methodological, architectural or computational resource differences.
This experimental setup allows for a clear evaluation of how well the federated setting bridges the gap between privacy preservation and segmentation accuracy, thus aligning with the core motivation of our study. To this end, the training process is conducted and empirically evaluated across four different setups:
@@
  • Local Training:  Each node independently trains its own nnU-Net model using only its local dataset. This setup serves as a baseline, reflecting the performance of a node/institution without participation in the federation.
@@
  • Centralized Training:   All data is aggregated into a single dataset, which is used to train a centralized nnU-Net model. This setup represents the theoretical upper bound of performance, though it is typically infeasible and often not recommended in real-world scenarios due to significant privacy risks and data-sharing constraints.
@@
  • Federated Fingerprint Training:   Utilizing our FFE approach (see Methods), the local dataset fingerprints from all nodes are aggregated at the central node into one federated fingerprint and distributed back to each node. As a result, each training site obtains an identical network architecture during the nnU-Net’s self-configuration process and enables the utilization of a standard FedAvg aggregation method. Moreover, this approach approximates nnU-Net’s behavior in the centralized training setup.
@@
  • Asymmetric Aggregation Training:   Implemented using our AsymFedAvg strategy, this allows nodes to maintain their original nnU-Net configured architectures while sharing compatible parts of their networks.
@@
For our experimentation, we maintain nnU-Net’s default configuration parameters as described in [8]. We test the 2D U-Net architecture on breast MRI and fetal ultrasound data and the 3D architecture on a cardiac MRI use case, consistently employing the default training length of 1000 epochs and using the last model checkpoint for evaluation across all federated nodes. In federated scenarios, all models are aggregated after each epoch and sent back to the clients. Following other recent works on medical image segmentation [8, 45, 46], we evaluate the models with Dice Similarity Coefficient (DSC) and 95% Hausdorff distance (HD95). For internal validation, we report results as the average of a 5-fold cross-validation, while for external testing, we follow the original nnU-Net approach by generating ensemble predictions from all five folds. Moreover, each experiment uses identical data splits—for example, the n-th fold in the centralized setup corresponds directly to the set of n-th folds from centers used in the local setup. To assess statistical significance between segmentation methods, we apply the two-tailed Wilcoxon signed-rank test on paired DSC scores with 5% significance threshold. By adhering to these protocols, we try to minimize the influence of other factors on the final results, ensuring that any observed performance variations are primarily due to our federated learning adaptations rather than modifications to the base framework.
Results
@@
Table 7: Quantitative results for multi-center MAMA-MIA and EuCanImage datasets. HD95 values reported in $m$. Best results for privacy-preserving methods are in bold.
@@
nnU-Net (local) 0.612 17.60 0.739 19.03 0.616 33.56 0.551 23.24 0.579 35.64 0.636 31.72 0.544 46.30
FednnU-Net AsymFedAvg 0.634 16.69 0.739 18.15 0.621 28.70 0.620 21.79 0.557 51.43 0.637 32.43 0.551 49.10
FednnU-Net FFE 0.644 16.20 0.742 17.48 0.653 39.54 0.660 17.49 0.582 28.83 0.645 33.52 0.565 40.41
nnU-Net (centralized) 0.654 17.28 0.750 19.84 0.657 58.99 0.675 17.35 0.582 52.44 0.644 41.17 0.569 57.98
Table 8: Quantitative results for multi-center M&Ms and ACDC datasets. HD95 values reported in $m$. Best results for privacy-preserving methods are in bold.
@@
nnU-Net (local) 0.908 4.27 0.912 3.62 0.902 6.56 0.896 4.31 0.900 4.34 0.887 4.79
FednnU-Net AsymFedAvg 0.909 4.20 0.913 3.70 0.907 4.23 0.897 4.29 0.902 4.07 0.889 4.58
FednnU-Net FFE 0.911 4.05 0.914 3.67 0.922 3.30 0.904 3.98 0.904 3.87 0.897 3.87
nnU-Net (centralized) 0.910 4.20 0.914 3.53 0.923 3.26 0.905 3.87 0.907 3.81 0.898 3.89
@@
Method Algeria Egypt Malawi Spain
external
Sierra Leone
@@
nnU-Net (local) 0.952 11.87 0.931 24.21 0.926 26.83 0.942 29.82 0.831 53.20
FednnU-Net AsymFedAvg 0.951 11.16 0.941 18.29 0.939 16.20 0.944 25.42 0.851 43.44
FednnU-Net FFE 0.955 10.23 0.954 14.56 0.943 15.64 0.938 47.58 0.809 46.57
nnU-Net (centralized) 0.954 10.23 0.953 14.50 0.943 15.55 0.932 55.76 0.804 54.34
Breast MRI
Table 7 summarizes segmentation performance on the MAMA-MIA dataset. Both FFE and AsymFedAvg outperform local training across all centers in DSC and HD95. FFE achieves average DSC improvements over local models on ISPY1 (3.2%, p=0.009), ISPY2 (0.3%, p=0.858), DUKE (3.7%, p=0.006), and NACT (10.9%, p¡0.001), also closely matching centralized results in DSC and surpassing them in HD95. Notably, AsymFedAvg yields the best HD95 on DUKE, reflecting its advantage in handling heterogeneous data (DUKE’s pixel size is 2× smaller; see Table 3). Fig. 4 highlights this heterogeneity, with ISPY2 models generalizing better to DUKE due to similar resolutions. Sample predictions (Fig. 5) show marked improvements for NACT and DUKE using federated methods. External validation on EuCanImage confirms FFE’s generalization, achieving DSC comparable to centralized training. KAUNO center results are consistent with MAMA-MIA, while GUMED’s drop likely reflects annotation discrepancies which are illustrated in Fig. 5.
Cardiac MRI
For M&Ms, federated methods outperform local models on both metrics, except for Clinica Sagrada Familia, where HD95 is slightly better locally. FFE consistently achieves the best DSC scores across internal and external datasets (combined significance using Fischer’s method: p¡0.001), while AsymFedAvg shows limited benefit—likely due to the dataset’s homogeneity, where unified architectures are more suitable. Cross-center results (Fig. 4) support this, with models generalizing well across unseen centers. Importantly, FFE matches centralized performance despite additional privacy constraints. Sample predictions (Fig. 5) show consistent results for central slices, with more variability at the apex (ACDC) and atria (Vall d’Hebron).
Fetal Ultrasound
In the fetal abdomen segmentation task, FFE and AsymFedAvg differ markedly (Table 9). Strong heterogeneity (capture devices, resolution; Table 6) impairs generalization from Algeria, Egypt, and Malawi to higher-resolution Spanish samples (Fig. 4). Unified architectures (FFE) are less optimal here: allowing Spain to retain its native architecture (AsymFedAvg) improves performance for external centers like Sierra Leone, notably the DSC raised to 0.851 (p ¡ 0.001). Fig. 5 shows that local models often misidentify abdomen regions, especially for Sierra Leone, whereas federated approaches improve delineation.
@@
Figure 5: Sample predictions of each method for datasets used in the study.
Discussion
Key Contributions and Performance
@@ Discussion
This study introduces FednnU-Net, a federated learning framework for 2D and 3D medical image segmentation that combines privacy preservation with high performance through two key innovations: Federated Fingerprint Extraction for standardized architecture sharing, and AsymFedAvg for flexible aggregation across heterogeneous node-specific models. Our experiments across three diverse medical imaging clinical tasks in (i) breast MRI (MAMA-MIA), (ii) cardiac MRI (M&Ms), and (iii) fetal ultrasound (AIMIX) - show that FednnU-Net reliably matches or surpasses the performance of centralized strategies across settings and modalities. These results add to a growing body of evidence illustrating that distributed model optimization can achieve comparable accuracy while simultaneously fulfilling the requirements of complex real-world clinical collaborations, including those related to data privacy and deviations in institution-specific protocols.
Towards Trustworthy AI in Healthcare
[…]
While some federated segmentation methods report high accuracy by tailoring solutions to narrow, task-specific settings, this work focuses on delivering an accessible, privacy-preserving framework that emphasizes ease of use and adaptability. By prioritizing broad applicability and real-world viability, the framework is positioned to support diverse research settings and facilitate widespread adoption.
Indeed, the ongoing shift towards privacy-preserving, decentralized deep learning solutions in healthcare is critical for protecting patient confidentiality [24]. In addition to this, the FUTURE-AI General 2 guideline states the necessity of implementing rigorous data protection throughout the AI lifecycle, including the use of privacy-enhancing technologies and secure data governance. In the context of FL with FednnU-Net, data protection is operationalized by ensuring that patient data never leaves its local institutions of provenance, reducing the risk of sensitive information leakage during collaborative model training. To this end, FednnU-Net addresses critical privacy and security concerns inherent in decentralized healthcare AI solutions, fostering trust among stakeholders and ensuring compliance with ethical and regulatory standards. While FednnU-Net addresses critical requirements for privacy and local adaptation in line with FUTURE-AI guidelines, several important aspects still warrant further investigation:
@@
    Scalability and Client Diversity: Future work is to assess FednnU-Net across an even larger, geographically dispersed federation with diverse patient populations [47] to study convergence dynamics at international, multi-institutional scales.
@@
    Heterogeneous Computational Resources: While our framework already supports diverse node capabilities, further systematic ablation studies under varied hardware conditions, including low-resource settings [48], are motivated towards showing the versatility for clinical adoption of the framework.
@@
    Enhanced Privacy Mechanisms: FL improves data privacy by design, but it still does not guarantee full security. Attacks such as model inversion [49] or membership inference [50] can still extract sensitive information from shared model updates. Integrating differential privacy [51, 52, 53, 54], synthetic data generation [55, 56, 57], and secure multiparty computation[58] could further reduce the risk of data leakage, even in aggregated updates.
Conclusion
FednnU-Net fills a critical gap by balancing segmentation performance, ease of use, and privacy preservation. As FL gains adoption in healthcare [24], our FednnU-Net framework is a key component in enabling the translation of high-performance deep learning into privacy-compliant, scalable clinical solutions. Our framework enables researchers and practitioners to easily integrate nnU-Net, along with other dedicated frameworks and innovations built upon it, such as for instance Bayesian uncertainty estimation [21], continual learning [18], and curriculum learning extensions [16], into FL setups, thereby broadening their applicability under privacy constraints and supporting the increasingly prevalent multi-center collaborations. Continued evaluation, targeted enhancements, and the combination with such innovations is to further strengthen the versatility and robustness of the framework, paving the way for widespread, trustworthy deployment across diverse clinical environments.
Data availability
Our framework’s code is publicly available at https://github.com/faildeny/FednnUNet. The MAMA-MIA dataset access information is available at https://github.com/LidiaGarrucho/MAMA-MIA. The access to the updated version of M&Ms1 The M&Ms2 datasets and annotations for fetal US will be provided through https://github.com/faildeny/FednnUNet upon the publication of this article. For any queries regarding the data used in this study, please contact the corresponding author.
Funding information
This work received funding from the European Union’s Horizon Europe research and innovation programme under Grant Agreement No. 101057849 (DataTools4Heart). This work has been supported by the European Union’s research and innovation programmes: Horizon Europe under Grant Agreement No. 101057699 (RadioVal) and Grant Agreement No. 101044779 (AIMIX), Horizon 2020 under Grant Agreement No. 952103 (EuCanImage).
Author contributions statement
The study was conceived by G.S. and designed by G.S., C.M.-I., and K.L. G.S. developed the methods, wrote the program code, and conducted the experimental work. F.A. and C.M.-I. contributed to code development and experimental work. F.A., D.A.K., R.O., and E.R.P. contributed to manuscript writing. G.S. led the writing of the manuscript. G.S., F.A., D.A.K., and R.O. revised the article. C.M.-I., F.A., L.G., D.A.K., and E.R.P. worked on data preparation. M.B., K.G., X.B., and P.J. provided the imaging data and annotations for this study. M.B., K.G., X.B., P.J., C.M.-I., E.R.P., O.D., K.K., and K.L. reviewed the manuscript. K.K. and K.L. supervised the study. K.L. provided funding for the study. All authors have read and approved the final manuscript.
Additional information
Competing interests The authors declare no competing interests.